You may want to reference (Pages 425-430) Section 10.5 while completing this problem. A 1.5-L, container of liquid nitrogen is kept in closet measuring 1.1 m by 1.2 m by 20 m ▼ Part A that the container is completely ful, that the temperature is 22.9 C, and that the atmospherk pressure is 12 atm, calculate the percent (by volume) of air that would be displaced if all of the liquid nitrogen evaporated (Liquid nitrogen has a density of 0 807 g/mL) Express your answer using two significant figures.
Answer #1

1.5 litres with density 0.807g/ml = 1210.5g N2
Molar mass N2 = 28.014g/mol
1,210.5g = 1210.5/28.014 = 43.21mol N2

Volume of N2 at 22.9°C and 1.2atm pressure
PV = nRT
1.2*V = 43.21*0.082057*295.9
V = 874.31 litres = 0.87431m³

Volume of closet = 1.1*1.2*2.0 = 2.64m³

% displced = 0.87431/2.64*100 = 33.12%